OpenStudy (anonymous):

y''-4y=2; y1=e^(-2x) Use the method of reduction of order to find a second solution y2(x) of the homogeneous equation.

OpenStudy (freckles):

\[y_2(x)=v(x)y_1(x)=v(x)e^{-2x}\] So we should have a solution of this form. So find y'_2 and y''_2

OpenStudy (freckles):

And we are going to plug that back in

OpenStudy (anonymous):

would it be y'_2 (x)=v'(x)e^(-2x)-2v(x)e^(-2x) y''_2(x)=v''(x)e^(-2x) -2v'(x)e^(-2x) -2v'(x)e^(-2x) +4v(x)e^(-2x)

OpenStudy (freckles):

ok and really just need that last on you found because we didn't have a y' in the equation

OpenStudy (freckles):

\[y''-4y=2 \\(v''e^{-2x}-4v'e^{-2x}+4ve^{-2x})-4(ve^{-2x})=2\]

OpenStudy (freckles):

that can be simplified just a little.

OpenStudy (freckles):

anyways... if we let v'=w then we have \[w'e^{-2x}-4we^{-2x}=2 \\ w'-4w=2e^{2x}\] Solve this linear differential equation then you can find v by integrating both sides of v'=w

OpenStudy (freckles):

one sec while i put a potatoe in the oven

OpenStudy (freckles):

back

OpenStudy (anonymous):

how would you solve the linear differential equation?

OpenStudy (freckles):

you could find the integrating factor

OpenStudy (freckles):

do you know how to solve diff equations in the form y'+p(x)y=q(x)

OpenStudy (anonymous):

i don't know what to do once i find the integrating factor.

OpenStudy (freckles):

let say the integrating factor you got was v \[(vy)'=2ve^{2x}\] integrate both sides

OpenStudy (anonymous):

the integrating value is e^(-2x^2) so it would be w'(e^(-2x^2)+4w(e^(-2x^2)=e^(-2x) x e^(-2x^3)